40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 160 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  8834   Fri Jul 12 09:00:10 2013 SteveUpdateGeneral Halogen illumination cables disconnected

Alex and Steve,

Old halogen chamber illuminator cabling disconnected and potenciometer board removed at 1Y1 in order to give room for pd calibration fibre set up.

Attachment 1: 1Y1oldilluminatersDisconnected.jpg
1Y1oldilluminatersDisconnected.jpg
  8833   Fri Jul 12 00:12:41 2013 KojiUpdateLSCPRMI + Y arm trial

[Koji, Manasa, Annalisa]

I made several trials to scan the arm on the IR TEM00 resonance while the PRMI was held with REFL165I&Q.
It was so hectic to keep multiple systems running correctly. We talked about how it should be automated.
We'll gradually offload the switching works on scripts.

In a good alignment condition, when I swept on the resonance, everytime the PRMI lost the lock. It reacquired
once the arm passed the resonance.

Lately I got difficulty to acquire lock of the PRMI while the arm is waiting at its off resonance.
If I change the ALS offset I got a stable lock in a certain offset, and did not get in another offset
so there could be something systematic. (The arm was in between the carrier resonance and the next sideband (55MHz) resonance).

-----

Procedure

[Preparation]

- Run LSCoffset script.

- Misalign PRM. Lock and align the arms with ASS.

- Go into the tables. Align the oplevs for ETMX/Y, ITMX/Y, and BS. (Very important for alignment stability)

- Align PRMI and lock PRMI. Unlock once.

- Go into the BS/PRM table. Align the oplev for PRM.

[ALS]

- Misalign PRM by -0.2

- Find the beat note at around 50MHz by changing the Yarm SLOW control. Today the PSL SLOW was ~0.24, and the Yarm SLOW was -10981.

- Reset Phase Tracker History (Important)

- Engage Yarm ALS with FM5. Tested the sign of the servo by giving 0.01 or -0.01. In my case, the negative number worked fine.
  Gradually increase the gain up to -10. Turn on FM2/3/6/7/10.

- Use Filter module "C1ALS-OFFSETTER2" to give the ALS sweep. I used FM1 (30mHz LPF). Change the offset while looking at the IR TRY and POY11 error signal.

- Once the resonance is found, shift the beat note by giving +10 or -10 offset.

[PRMI]

- While the arm is kept off resonance, align PRM.

- Lock PRMI with REFL33I and AS55Q. Turn on PRM ASC.

- Once the stable lock is obtained, switch the input signals to REFL165I&Q. I used REF33I x1.0->REFL165I x0.8 and AS55Q x1.0 -> REFL165Q x0.5

[PRMI + one arm]

- Revert the ALS offset by 10 to bring the arm on the resonance the see what happens.

 

  8832   Thu Jul 11 23:50:57 2013 gautamConfiguration PZT Driver Board-changes made

 Summary:

Continued with tests on the PZT driver board. I made a few changes to replace defective components and also to modify the gain of the HV amplifier stage. I believe the board has been verified to be satisfactory, and is now ready for a piezo to be connected, tested and calibrated.

Changes made:

  • I tested the board with the full 100V bias voltage today, working my way up from 30V in steps of about 20V and verifying the output at each stage.
  • In order to deliver 100V to the board, it was necessary to change the maximum current limit on the KEPCO supply, which is set at default at ~1.6 mA. The KEPCO power supply placed near rack 1X2 (which I believe was used to power a piezo driver board) is labelled 150V, 12 mA, though I found that the board only drew 7mA of current when the power supply output 100V. I have set the limit to 10 mA for the time being.
  • The potentiometer in the third stage (R44 in the schematic) was faulty so I replaced it with another 100K potentiometer, which was verified to work satisfactorily.
  • We expect the DAC output to supply a voltage to the input of the PZT driver board in the range -10V to 10V. Today, I verified this by using my temporary break-out cable. I hooked this up to the DAC at 1Y4 and output a 3 Hz sine wave with amplitude of 32000 counts (the maximum) on channel 9. The output as observed on an oscilloscope (image attached) was a 10Vpp sinusoid, confirming the above hypothesis. As mentioned in my previous elog, the gain of the high-voltage amplifier stage is ~15, which would mean the output would saturate if the input were to be >6V. I have changed the gain of all 4 stages (M1-pitch, M1-yaw, M2-pitch and M2-yaw) to ~4.85 by swapping the 158k resistors (R43, R44, R69 and R70 in the schematic) for 51k resistors. 
  • It was necessary to change the value of the biasing potentiometers after the change in gain so that 0 input voltage once again provided 50V at the output, as required by the PZTs for there to be no tilt. This was done and verified. This biasing voltage now is ~10.4V in all four stages.
  • Having adjusted the gain, I tested the circuit over the expected full range of the input voltage from the DAC (from -10V to 10V) from the DS345 function generator (0.05Hz sinusoid). I monitored the output using a multimeter, as the monitor channels were peaking at ~7V, which was above the limit for the oscilloscope I was using. It was verified for all four channels that the output was between 0 V and 100 V (the safe range quoted in the datasheet for the tip-tilts, for this range of input voltages. So I think we are ready to connect a PZT to the board and conduct further tests, and calibrate the PZT. 

Pending Issues:

  • Koji pointed out that there has to be an anti-imaging filter stage between the DAC output and the filter stage, which I had not considered till this point.Another subtle point is that the DAC output is differential while the driver boards have a single-ended input, which means we effectively lose half the range of the PZTs. 
  • A suitable candidate is the D000186-rev D. Some information about the present state of this board is detailed in this elog. This board also solves the problem of the differential vs single input as the input to the AI board is differential while the output is single-ended. Koji has given me one of the boards he had collected. 
  • Some changes will have to be made to this version of the board in order to make it compatible with the existing DAC. I will first have to measure the power spectrum of the DAC output to verify that the AI boards need notches at 64k and 128k. The existing notches are at 16k and 32k, and once the DAC power spectrum has been verified, I hope to affect the necessary changes by switching out the appropriate capacitors on the existing board. 
  • The AI board is an extra element which I have now added to an updated wiring diagram, attached.

Revised Wiring Diagram:

ASC_schematic.pdf

 

DAC Max. Output Trace on Oscilloscope

 

DAC_Max_output.JPG

 

 

 

  8831   Thu Jul 11 14:38:38 2013 AnnalisaUpdateGreen LockingY arm cavity scan

Yesterday I did a cavity scan with IR while holding the Yarm with green.

ALS servo tuning:

  • C1ALS-BEATY_FINE_PHASE

             The gain of the loop is set such that BEATY_FINE_Q_ERR x GAIN = 120k. This is a kind of "empirical low" in order to have the UGF around 1kHz. 

  • C1ALS_YARM

             Start with FM5 [1000:1] enabled, determine the sign of the gain increasing it in small steps and making sure that the mirror doesn't get a kick. Then gradually raise it while looking at the BEATY_PHASE_OUT power spectrum.

             Enable FM7 [RG16.5], FM6 [RG3.2], FM3 [1:5], FM2[0:1], FM10 [40:7].

Plot 1 shows the power spectrum of BEATY_PHASE_OUT (calibrated in Hz).

  1. blue curve - ALS disabled
  2. green curve - in loop measurement, ALS enabled and servo tuned as described above
  3. grey curve - RMS of the in loop measurement
  4. red curve - out of loop measurement (arm locked with IR)
  5. pink curve -  RMS of the out of loop measurement

Offset setting and cavity scan

The C1ALS_OFFSETTER2 was used to set an offset for ALS scan.

  • LPF30m enabled
  • Ramp time set to 150s
  • Offset set to 1500 (approximately 3 FSR in this interval)

Many scans have been done to find the optimal offset conditions, I only attached one (Plot 2).

I also misaligned the END mirror in pitch to enhance the HOMs peaks, but it turned out that it was not enough, because I didn't see a very big difference between the "aligned" and the "slightly misaligned" measurements (Plot 3). 

NEXT STEPS

Increase the cavity misalignment both in pitch and in yaw and repeat the measurement.

 

Attachment 1: IRlocked_res_mot_0710.pdf
IRlocked_res_mot_0710.pdf
Attachment 2: ArmScanN5.png
ArmScanN5.png
Attachment 3: CavityScanETMYmis
  8830   Thu Jul 11 13:52:51 2013 CharlesUpdateISSRMS threshold detection and triggering

There are essentially two major portions of the ISS I am designing. One system has the voltage reference, differential amplifier and filtering servo (schematic attached) while the other has a comparator circuit and a triggering mechanism. The first system amplifies an error signal obtained from the PD output and the voltage reference, which is then fed back through the AOM. I've done a lot of work designing/prototyping this first half and now I'm starting to design the second half.

The second system's main purpose is to maintain loop stability as the ISS is engaged. Let's assume a user has decided they want noise suppression. They would first close the ISS feedback loop and an error signal would pass through three unity-gain buffers, providing minimal noise reduction. The user can then send a signal to theTRIGGER 1 port to switch the first stage from its unity-gain position to its filtering position and reduce the intensity noise further. This signal will most likely be digital in origin. Alternatively, when the user first closes the ISS loop, the first stage can already be in its filtering position rather than necessitating two commands.

A test channel (not drawn in the included schematic) will monitor the RMS level of the incoming signal from the PD. This noisy AC signal will first be amplified and then passed through an RMS-to-DC converter. The resulting DC signal is used as a part of the triggering mechanism for later stages. Once the first stage has been switched manually, and the DC signal corresponding to RMS noise of the PD output drops below a certain threshold, stages 2 and 3 will be internally triggered with a short delay between them. Toward being able to detect this threshold, I have designed a simple comparator circuit with an LT1016. The circuit has a fairly low-level output when the input voltage is larger than the threshold (about 1.6 V for my simple prototype), but when the input passes below the threshold, the comparator puts out almost 4 V, a number limited by the supply voltage. The schematic is shown below.

Simple_Comparator_Circuit.png

The component V2 and the various voltage dividers serve to establish the reference/threshold voltage. Note that although the LT1016 is not powered in the schematic, it requires ±5 V (a max of 7 V). The above circuit was also prototyped on a breadboard and I characterized it with an oscilloscope. Using a CFG253, I made a low frequency (~0.3 Hz) triangle wave with an amplitude and DC offset such that it oscillates between 0 and 5 V. This was applied to the IN node in the above schematic. The input waveform and the circuit's response (voltage at the OUT node) are shown below. As expected, R2 serves to establish hysteresis. The comparator switches to 'high' output until the input drops below 1.6 V, and then it doesn't switch back to the 'low' output until the input goes up to ~3.4 V.

F0001TEK.JPG

This behavior is ideal for our application as we can detect when the DC signal from the RMS-to-DC converter drops below a certain level (i.e. the first stage that has been activated does some amount of filtering to lower RMS noise), and then we can trigger subsequent filter stages off of the comparators high-level output. 

This circuit could easily be used to drive the MAX333a switches shown in the first schematic attached. I believe the low-level output is not sufficient to switch the MAX333a although the ~4 V high-level output is quite sufficient. Regardless, these exact values (thresholds, outputs etc) will be determined after I have a better idea of the RMS noise of the laser without any intensity stabilization as well as a solid understanding of how the AD8436 RMS-to-DC converter works. This was simply a proof of concept for lower threshold detection using basic Schmitt trigger topology.

Attachment 1: 40mServo_v1.pdf
40mServo_v1.pdf
  8829   Thu Jul 11 12:00:50 2013 AlexUpdate Planned rack additions

[Eric, Alex]

We mounted our Laser Module and Laser Power Source in rack 1y1. We plan to add our RF Switch and Transformer Module to the rack, as pictured. (Note: drawn-in boxes in picture are approximately to scale.) Note that the panel of knobs which the gray boxes overlap is obsolete and will soon be removed.

Attachment 1: Annotated_Rack_1y1_-_update.pdf
Annotated_Rack_1y1_-_update.pdf
  8828   Thu Jul 11 09:24:03 2013 SteveUpdateGeneralthere should be nothing left on optable tops

There are 4 oscilloscopes left on the AP optical table top.... It's only 25 lbs... Do not leave anything on the optical table tops!

  8827   Thu Jul 11 09:15:10 2013 SteveUpdateendtable upgradeETMY optable grounded

ETMY optical table top was grounded to the ETMY chamber through 1 Mohms this morning. I  also strain releifed relieved a few cables that were pulling on components directly.

  8826   Thu Jul 11 07:34:42 2013 manasaUpdateLSCYarm held nicely on IR resonance with ALS, PRMI+arm attempt

Quote:

We knew that the Yarm was well aligned, since our IR resonance was > 0.98, but it had been a while since we had aligned the X arm. 

 The X arm was locked with TRX>0.98 earlier last night while I was measuring the out of loop noise of the phase tracker.

  8825   Thu Jul 11 03:14:19 2013 JenneUpdateLSCYarm held nicely on IR resonance with ALS, PRMI+arm attempt

[Annalisa, Jenne, Nic]

After having troubles with the Xarm earlier (maybe Manasa can write/say something about this?  Something about perhaps seeing the phase tracker jump, and cause it to lose lock?), we moved on to the Y arm. 

Annalisa locked the Yarm green, and closed the ALS loop.  I believe that earlier today, she tuned the gain such that we don't start getting gain peaking at a few hundred Hz.  We would like to get a script going, so that it's not so labor intensive to reclose the ALS loop after an MC lockloss....but that's a daytime task.

We then found the IR resonance, using only the Yarm ALS system.  After Manasa's work yesterday, the Yarm was very stable while locked with the ALS.  We took a power spectrum of POY11_I_ERR, which I have calibrated using the number in elog 6834 of 1.4e12 cts/m, or 7.14e-13 m/ct.  See the figure below.

After that, we changed the offsetter2 offset such that the arm was off resonance, but not so far off that we crossed any significant resonances (in particular, we wanted to not go as far as the 55MHz resonance). 

Then, I tried to lock the PRMI for a while, but the alignment wasn't very good.  We knew that the Yarm was well aligned, since our IR resonance was > 0.98, but it had been a while since we had aligned the X arm.  I tweaked the ITMX position to make the Michelson dark, and then tried acquiring PRMI lock.  At first, I tried with REFL165 I and Q, but with the non-ideal alignment and the offset in the 165 diode (LSC offsets was not run this evening), I wasn't catching any locks.  I then switched to AS55Q and REFL33I, but wasn't able to catch lock there either. 

The MC lost lock, which made us lose the ALS loop, but the ALS had been locked for more than 30 minutes, at least.  I tried locking the PRMI with the current alignment (after having misaligned ETMY), but was only able to get lock stretches of 1 second at maximum.

We are calling it a great success for the night, since we have confirmed that, at least for the Yarm, Manasa's beatbox work has improved things.  Also, we have a pretty solid plan for trying the PRMI+arm tomorrow, even though it didn't work out tonight.

Attachment 1: Yarm_onIRresonance_noPRMIyet_POYcalibrated.pdf
Yarm_onIRresonance_noPRMIyet_POYcalibrated.pdf
  8824   Thu Jul 11 00:30:27 2013 manasaUpdateGreen LockingX arm ALS post-beatbox makeover

I ran a series of diagnostics on the X arm ALS to look at how the beatbox behaves after the makeover.

Diagnostic tests run:
1. X arm ALS in-loop spectrum
2. X arm ALS out-of loop spectrum
3. X ALS scan of the X arm cavity

The noise suppression looks better after the makeover at the lower frequencies. To suppress the noise at high frequencies, we would have to add more whitening filters.

Attachment 1: XALS_inloop.pdf
XALS_inloop.pdf
Attachment 2: XALS_scan.pdf
XALS_scan.pdf
Attachment 3: ALS_outloop.pdf
ALS_outloop.pdf
  8823   Wed Jul 10 22:41:06 2013 gautamConfigurationendtable upgradePZT Driver Board

 I did the following with the PZT Driver Board: 

 

  •  With an expansion card attached to the driver board, I used an Agilent E3620A power supply to verify that the 15V and 24V supplies were reaching the intended ICs. It turns out that the +24 V supply was only meant to power some sort of on-board high voltage supply which provided the 100V bias for the PZTs and the MJE15030s. This device does not exist on the board I am using, jumper wires have been hooked up to an SMA connector on the front panel that directly provides 100V from the KEPCO high voltage supply to the appropriate points on the circuit.

  •  All the AD797s as well as the LT1125CS ICs on the board were receiving the required +15V.

      

The next step was to check the board with the high-voltage power supply connected.

 

  •  The output from the power supply is drawn from the rear output terminal strip of the power supply via pins TB1-2 (-OUT) and TB1-7 (+OUT). I used a length of RG58 coaxial cable from the lab and crimped a BNC connector on one end, and stripped the other to attach it to the above pins.

  •  There are several options that can be configured for the power supply. I have left it at the factory default: Local sensing (i.e. operating the power supply using the keypad on the front of it as opposed to remotely), grounding network connected (the outputs of the power supply are floating), slow mode, output isolated from ground.

  • I was unsure of whether the grounding network configuration or the 'positive output, negative terminal grounded' configuration was more appropriate. Koji confirmed that the former was to be used so as to avoid ground loops. When installed eventually, the eurocrate will provide the ground for the entire system.
  • I then verified the output of the HV power supply using a multimeter from 2V up to 150V.
  • I then connected the high voltage supply to the PZT driver board with a BNC-SMA adaptor, set, for a start, to output 30V. Ensured that the appropriate points on the circuit were supplied with 30V.

 

I then hooked up a function generator in order to simulate a control signal from the DAC. The signal was applied to pin 2 of the jumpers marked JP1 through JP4 on the schematic, one at a time. The signal applied was a 0.2 Vpp, 0.1 Hz sine wave.

 

 

 

  •  The output voltage was monitored both using a DMM at the SMB output terminals, and at the monitor channels using an oscilloscope. The outputs at both these points were as expected.
  • There are 4 potentiometers on the board, which need to be tuned such that the control output to the piezos are 50V when the input signal is zero (as this corresponds to no tilt). The gain of the amplifier stage (highlighted in the attached figure) right now is ~15, and I was using 30V in place of 100V, so an input signal of 2V would result in the output saturating. This part of the circuit will have to be tuned once again after applying the full 100V bias voltage. 
  • Koji suggested decreasing the gain of the amplifier stage by switching out resistor R43 (and corresponding resistor in the other 3 stages on the board) after checking the output range of the DAC so that possibility of unwanted saturation is minimised. I need to check this and will change the resistors after confirming the DAC output range. 
  • The potentiometers will have to be tuned after the gain has been adjusted, and with 100V from the high-voltage DC power supply. 

  

To Do:

 

  • Switch out resistors
  • Tune potentiometers with 100V from the HV supply
  • Verify that the output from the board after all the tuning lies in the range 0-100V for all possible input voltages from the DAC.
  • Once the output voltage range has been verified, the next step would be to connect a PZT to the board output, affix a mirror to the tip/tilt, and perform some sort of calibration for the PZT. 

HV_Amplifier.pdf

 

 

 

 

 

 

  8822   Wed Jul 10 14:49:00 2013 SteveUpdateGreen Locking more fiber protective tubing layed

Quote:

Craig, Gautam and Steve,

Single mode fiber 50m long is layed out into cable tray that is attached to the beam tube of the Y arm.

It goes from ETMY to PSL enclosure. It is protected at both ends with " clear- nylon slit corrugated loom tubing " 1.5" ID

The fiber is not protected between 1Y1 and 1Y4

 Installed 0.5" ID 10 ft long protective tubing at the PSL end of the  ETMY fiber this morning. Here I had to cable tie a bunch of cables at the east side of the PSL enclosure.

They were hanging off the table blocking space were the sliding doors move.

 At the ETMX end of the X-arm fiber received the same protective tubing.

Attachment 1: fibreETMXprotectedendOnly.jpg
fibreETMXprotectedendOnly.jpg
  8821   Wed Jul 10 11:44:02 2013 AnnalisaUpdateGreen LockingY beat note found!

I found the beat note for the Y arm. Nothing was changed with respect to yesterday night, but the beat is back!

  8820   Wed Jul 10 11:27:02 2013 manasaUpdateGreen LockingX arm beatnote found

I found the beat note for X arm. I did not change anything this morning (to the best of my knowledge). Hooking up the spectrum analyzer, I could find the beatnote signal at the PD RF output, after the amplifier and also at the MON port of the beatbox. I still don't know what changed from the last night set of trials

  8819   Wed Jul 10 02:28:04 2013 AnnalisaUpdateGreen LockingBeat notes lost!

[Manasa, Jenne, Annalisa]

I was going to find the beat note to start the cavity scan, but I couldn't.

These are the steps I followed:

  • locked the arm with IR to reduce the arm swinging
  • locked the green on the arm
  • started changing the green temperature setting the offset from the slow servo2 in the ALS. The PSL slow actuator ADJ was always set approximately to zero, and the PSL temperature was checked in order to set the auxiliary laser temperature where the beat was expected (as in the plot)

After spanning the temperature by approximately 4degC, we started be suspicious that I couldn't find the beat in the range of temperature where it was supposed to be found, and we started making several trials:

  • PD output disconnected from the beatbox and connected to the cable running to the Control Room
  • Checked that the cable going to the Control Room was working by sending a signal with the Marconi (the cable was working)
  • Put back the amplifier that had been previously removed
  • PD DC output checked with the oscilloscope
  • Spectrum analyzer connected to the PD output without passing trough the cable

The same trials were done also for the X arm, but we didn't succeed in finding the beat for the X neither.

 

  8818   Wed Jul 10 02:10:41 2013 manasaUpdateGreen LockingBeatbox gets a makeover

Quote:

[Koji, Manasa]

I wanted to investigate on the ALS electronics(in particular the beatbox and the phase tracker) and find out if the beatbox is showing a linear behavior
as we expect it to and as to why we have been seeing sudden jumps at the phase tracker output.

I have been using the Xarm part of the beabox.
I used Marconi as well as signal generator to do frequency sweep/modulation at the RF input of the beatbox and looked at the I_MON output of the beatbox.

We observed sudden jumps in the beatbox output from time to time while we either varied the carrier frequency or the RF amplitude.
Also the beatbox output shows high frequency oscillations at ~95MHz (source unknown). It is for sure that the beatbox is not behaving the way it should
but we could not tell more or troubleshoot with the beatbox mounted on the rack.

I am going to let Annalisa do her Y arm ALS scan tonight and pull out the beatbox tomorrow to fix it.

 The beatbox output showed high frequency oscillations during the troubleshooting process yesterday. I removed the beatbox from the rack. With no RF inputs, just powering the beatbox showed these high frequency oscillations at the beatbox output. This confirms that these oscillations are from the op-amp AD829JR. I replaced these with low noise OP27G. Also I removed the AD829JR that were soldered to the frequency divider and comparator which are not being used. Output buffer U10 was also removed.

After replacing with OP27G, I rechecked the beatbox with and without the RF input. There were no more high frequency contaminations and beatbox seemed to behave as it is supposed to when a frequency modulated RF input is fed. I put the beatbox back on the rack and did  a quick recheck.

Before (top) and after (bottom) pictures

IMG_0842.JPGIMG_0844.JPG

IMG_0845.JPGIMG_0846.JPG

 

  8817   Wed Jul 10 01:27:44 2013 gautamUpdateGreen LockingY-end Green PDH open-loop transfer function

 [Annalisa, gautam]

Summary:

We have measured the open-loop transfer function of the Y-end green PDH loop. From the measurement, the loop UGF is ~12kHz.

Details:

We have been trying to measure this transfer function for some time now, and playing around with various points of injecting the excitation and measuring the output. Koji helped arrive at one that actually worked, and the scheme used to make this measurement is shown in the sketch below. The SR785 signal analyzer was used to make the measurement, while an SR560 preamp was used to sum the output from the PDH box (PZT-OUT) and the excitation, with this sum being delivered to the auxiliary laser PZT via a pomona box that sums the servo output and the signal from the LO. The transfer function measurement made was a1/a2 w.r.t the sketch attached.

  • The swept-sine measurement was done from high to low frequencies, as the open-loop gain was expected to be high at low frequencies.
  • After some trial and error, we realised that the excitation amplitude on the SR785 can be varied continuously during the course of a swept sine measurement using the dial on the front panel. We started out with a 1mVpp signal at the high end of the frequency sweep (~102kHz, the upper limit on the SR785) and went up to 17mVpp at ~30Hz). These values were determined by trial and error, and were approximately the maximum that did not kick the loop out of lock/into a higher order mode.

Remarks:

  • As per this paper, the expected bandwidth of this loop is expected to be ~30kHz, while the measured UGF was more like 11.7kHz. Perhaps we can get this closer to the expected 30kHz by increasing the servo gain. The measurement shown was done with the servo gain knob on the Universal PDH box set to ~7.86. We tried two other values, ~8.2 and 10 (this was the limit on the knob), but the UGF first increased to ~13kHz (for the 8.2 gain), and then decreased to ~5kHz with a gain of 10. Not sure why this was, but it can be looked into further. 
     

Set-up to measure Y-end Green PDH transfer function:

Green_PDH_measurement.pdf

 

Measured Open Loop Transfer Function:

Y-end_Green_PDH.pdf

  8816   Tue Jul 9 23:27:17 2013 KojiSummaryLSCMICH: ITMX/Y <=> PRM/BS

The MICH actuation with PRM/BS was investigated again.

(ITMX -1 / ITMY +1) is equivalent to (PRM -0.267 and BS +0.50).


- PRMIsb was locked with REFL33I&AS55Q.

- Using the locking module in the LSC model, actuate ITMX (-1) and ITMY (+1) at 580.1Hz. Note that the notch filters in the MICH/PRCL servos were on.

- Look at the peak in the AS55Q spectrum. Tune the BS element in the output matrix of the lock-in to minimize the peak height.
=> The peak was minimized at BS = -0.50.

- Look at the peak in the REFL33I spectrum. Tune the PRM element in the output matrix of the lock-in to minimize the peak height.
=> The peak was minimized at PRM = +0.267

- These measurement leads to the conclusion mentioned above.

  8815   Tue Jul 9 20:09:53 2013 KojiUpdateIOOWFS debugging

The low UGFs of the MC WFS servos made the MC insane thesedays:
The servos are too slow and we kept having significant misalignment left uncompensated.

I increased the total gain of the MC WFS from 0.01 to 0.4 (x40) to make the UGFs of the
WFS paths to ~2Hz. This was too much gain for the QPD path so the gains for the QPD paths
were reduced by a factor of 4 (x10 in total).

The script mcwfsup was also modified accordingly.

  8814   Tue Jul 9 18:44:37 2013 gautamUpdateCDSDAC Interface Board-Pin Outs

  Summary:

The pin-outs for the DAC interface board have been determined.

Details:

  • I used a temporary break-out cable (pic attached) and connected the 40pin IDC connector on this to the DAC interface board at 1Y4.
  • I had a hypothetical pin-out map which was to be verified. So I connected pairs of ribbon wire to an oscilloscope in the configuration which I believed to be correct, and then used awggui to send a 3Hz, 10000 count sine-wave to the corresponding channel via the excitation points set up earlier.
  • I verified that the correct waveform showed up on the scope screen. I then tried sending the same signal to another DAC channel and verified that there were no accidental shorts/bad connections. The signal was fairly noisy, but this was probably because of the makeshift connections.
  • Repeated the above for all 8 channels in the bank marked 9-16 on the DAC interface board.

Turns out that my deductions using the D0902496 wiring diagram, a spare D080303 DAC to IDC adaptor and a multimeter were correct! The pin outs as determined by this test are sketched in the graphic below.

To Do:

  •  Now that the pin-outs have been determined, I need to go about making the custom ribbon that will connect the 40pin IDC on the DAC interface board to the 10-pin IDC on the PZT driver board. Because there is a pair of wires that will have to be 'skipped' while going from the 40-pin to the 10 pin IDC (corresponding to the pair not-connected between two DAC channels on the 40-pin IDC), this may be tricky.

Misc:

The excitation points added to the simulink model are still there, I plan on keeping it as such till I finish installation of the boards as they will be useful for testing purposes.

 

Pin-Outs of the DAC to IDC Adaptor (D080303) inside the "DAC Interface Box at 1Y4":

DAC_Interface_Board_Pin-out.pdf

 

Makeshift break-out ribbon cable:

 

break-out_ribbon.JPG

 

 

  8813   Tue Jul 9 17:03:06 2013 SteveUpdateGreen Lockingfiber layed for Y arm

Alex, Gautam and Steve,

Single mode fiber 50m long is layed out into cable tray that is attached to the beam tube of the Y arm.

It goes from ETMY to PSL enclosure. It is protected at both ends with " clear- pvc, slit corrugated loom tubing " 1.5" ID

The fiber is not protected between 1Y1 and 1Y4

Attachment 1: fromETMYtowardPSL.jpg
fromETMYtowardPSL.jpg
Attachment 2: fibreETMYtoPSL50m.jpg
fibreETMYtoPSL50m.jpg
Attachment 3: PSLfiberfromETMY.jpg
PSLfiberfromETMY.jpg
  8812   Tue Jul 9 16:08:32 2013 JenneUpdateASCPOP QPD calibration attempt

[Jenne, Alex] 

Calibration data for the POP QPD has been taken, with the PRMI locked on sideband (with AS55Q and REFL33I, since it stayed locked longer with those 2).  ASC was on, AC coupled. 

We didn't get too far on either side of center of the QPD, since the ASC servo would go unstable, so we only explored the roughly linear region.  Data / plots / analysis to follow.

  8811   Tue Jul 9 12:01:20 2013 gautamUpdateCDSset up for testing DAC Interface-board pin outs

 

 Jenne just rebooted c1scy and daqd on the framebuilder. We will do the actual test after lunch.

  8810   Tue Jul 9 11:41:22 2013 JenneUpdateASCPOP QPD calibration attempt

I was bad, and forgot to elog the most important part of my work yesterday - that I had rotated the POP QPD by 90 degrees, so that I could fit the micrometer onto the table.  There is a sticker on the front of the QPD to indicate which direction is "X" and "Y" for the output of the readout box.  Right now (and the way that I will mount the QPD to the table, after I redo the calibration today), X is PITCH, and Y is YAW.  Koji and Nic swapped the cables to the ADC to make this all consistent.

Yesterday, I locked the PRM-ITMY half cavity, and tried to take calibration data.  However, with no ASC servo engaged, the beam was still moving.  Also, with only the half-cavity, I had very little light on the QPD, and since it has internal normalization, the outputs can get a little funny if there isn't enough light.  I had checked, and even with the gain cranked up to maximum, the "light level too low" LED was illuminated.  So, my calibration data from yesterday isn't really useful.

Today, hopefully after lunch, I will lock the PRMI with the new AC-coupled ASC servo, so that I can have the servo on, and the PRMI locked on the sideband, so that I have more light on the QPD. 

After that, it seems that the final thing we need to do before we vent is hold an arm near, but off resonance, lock the PRMI, and then swing the arm in and out of resonance a bit.

  8809   Tue Jul 9 11:37:37 2013 gautamUpdateCDSset up for testing DAC Interface-board pin outs

The bank marked channel 9-16 is free, but the connector is a 40 pin IDC and I need to know the exact pin-out configuration before I can set about making the custom ribbon cable that will send the control signals from the DAC card to the PZT driver board. 

The DAC interface board on rack 1Y4 seems to be one of the first versions of this board, and has no DCC number anywhere on it. Identical modules on other racks have the DCC number D080303, but this document does not exist and there does not seem to be any additional documentation anywhere. The best thing I could find was the circuit diagram for the ADL General Standards 16-bit DAC Adapter Board, which has what looks like the pin-out for the 68 pin SCSI connector on the DAC Interface board. Koji gave me an unused board with the same part number (D080303) and I used a multimeter and continuity checking to make a map between DAC channels, and the 40 pin IDC connector on the board, but this needs to be verified (I don't even know if what is sitting inside the box on 1Y4 is the same D080303 board).

Jenne suggested making a break-out cable to verify the pin-outs, which I did with a 40-pin IDC connector and a bit of ribbon wire. The other end of the ribbon wire has been stripped so that we can use some clip-on probes and an oscilloscope to verify the pin-outs by sending a signal to DAC channels 9 through 16 one at a time. On the software side, Jenne did the following:

  • Restarted the mx_stream on c1iscey  (unrelated to this work)
  • 8 Excitation points added in the simulink model on c1scy 
  • Model compiled and installed

We have not restarted c1scy yet as Annalisa is working on some Y-arm stuff right now. We will restart c1scy and use awggui to perform the test once she is done.

 Pink edits by JCD

  8808   Tue Jul 9 01:18:48 2013 Nic, KojiUpdateASCPRMI locking / PRM ASC adjustment

[Koji, Nic]

- Locked PRMI with REFL165 I/Q

- Aligned the POP beam on the QPD. We found that the vertical motion of the beam appeared in the yaw signal, and horizontal motion in the pitch signal.
  This was fixed by swapping the cables to the ADC. Later it turned out that this was caused by the calibration setup for the QPD.
  We requested Jenne to fix the QPD on the table with the current orientation.

- Re-implemented the AC-coupled ASC servo. The filters were just copied from the previous PRM ASC servo (in the SUS ASC filter).
  The same filter was installed to the pitch and yaw filter modules for now. The gains were adjusted to have some stable lock stretches.
  C1:ASC-PRCL_YAW_GAIN: -0.01
  C1:ASC-PRCL_PIT_GAIN: -0.01

  The power spectra of C1:ASC-PRCL_YAW_IN1 and C1:ASC-PRCL_PIT_IN1 were attached.
  The reference curves are the ones with the servo on. The other two are the free-running stability of the QPD output.

- Modified the up and down scripts for the PRM ASC for the new setup.
  It first turns on the inputs of the filters and then turn on FM2/3.
  It assumes that the outputs are engaged all time.

 

Attachment 1: PRMI_ASC.pdf
PRMI_ASC.pdf
  8807   Mon Jul 8 21:46:31 2013 manasaUpdateGreen LockingBeatbox

[Koji, Manasa]

I wanted to investigate on the ALS electronics(in particular the beatbox and the phase tracker) and find out if the beatbox is showing a linear behavior
as we expect it to and as to why we have been seeing sudden jumps at the phase tracker output.

I have been using the Xarm part of the beabox.
I used Marconi as well as signal generator to do frequency sweep/modulation at the RF input of the beatbox and looked at the I_MON output of the beatbox.

We observed sudden jumps in the beatbox output from time to time while we either varied the carrier frequency or the RF amplitude.
Also the beatbox output shows high frequency oscillations at ~95MHz (source unknown). It is for sure that the beatbox is not behaving the way it should
but we could not tell more or troubleshoot with the beatbox mounted on the rack.

I am going to let Annalisa do her Y arm ALS scan tonight and pull out the beatbox tomorrow to fix it.

  8806   Mon Jul 8 16:27:49 2013 AlexUpdate Planned rack additions

Alex and Eric

For the photodetector frequency response automation project, we plan to add modules to rack 1y1 as shown in the attached picture (Note: boxes are approximately to scale). 

The RF switch will choose which photodetector's output is sent to the Agilent 4395A Network Analyzer.

The Diode Laser Module is powered by Laser Power Supply, will be modulated by the Network Analyzer and will be output to a 1x16 optical splitter which is already mounted in another rack (not pictured). 

The Transformer Module has not been built yet.

We would like to install the power supply and the laser module tomorrow and will not begin routing fibers and cables until we post a drawing in the elog.

Also, our reference photoreceiver arrived today.

 

Attachment 1: Annotated_Rack_1y1.pdf
Annotated_Rack_1y1.pdf
  8805   Mon Jul 8 15:31:48 2013 JenneUpdateASCPOP QPD calibration prep

I am prepping to do the POP QPD calibration, and so have turned off the POP QPD, and put it onto a micrometer stage.  My plan is to (after fixing the ASC servo filters to make the servo AC coupled, rather than DC coupled) lock the PRM-ITMY half cavity, and use that beam to calibrate the QPD.  While this isn't as great as the full PRMI, the PRMI beam moves too much to be useful, unless the ASC servo is engaged.

While on the table, I noticed 2 things:

* In order to place the micrometer, I had to temporarily move the POP55 RFPD (which has not been used in quite a long time).  I think it's just that the panel-mount SMA connector isn't tight to the panel inside, but the RF out SMA cable connector is very loose.  I have moved the POP55 RFPD to the very very south end of the SP table, until someone has time to have a quick look. (I don't want to get too distracted from my current mission, since we haven't put beam onto that PD for at least a year).

* The ITMX oplev beam setup isn't so great.  The last steering mirror before the beam is launched into the vacuum is close to clipping (in yaw... pitch is totally fine), and the steering mirror outside of vacuum to put the beam on the QPD is totally clipping.  The beam is falling off the bottom of this last steering mirror.  Assuming the beam height is okay on all of the input optics and the in-vac table, we need to lower the last steering mirror before the oplev QPD.  My current hypothesis is that by switching which in-vac steering mirror we are using (see Gautam's elog 8758) the new setup has the beam pointing downward a bit.  If the problem is one of the in-vac mirrors, we can't do anything about it until the vent, so for now we can just lower the out of vac mirror.  We should put it back to normal height and fix the oplev setup when we're at atmosphere.

  8804   Mon Jul 8 13:45:19 2013 gautamConfigurationendtable upgradeDriver board verification

With the help of an expansion card,  I verified that the + 15V and + 24V from the eurocrate in the slot I've identified for the PZT driver boards are making their way to the board. The slot is at the right-most end of the eurocrate in 1Y4, and the rack door was getting in the way of directly measuring these voltages once I hooked up the driver board to the expansion card. So I just made sure that all the LEDs on the expansion card lit up (indicating that the eurocrate is supplying + 5, + 15 and + 24V), and then used a multimeter to check continuity between the expansion card and the driver board outside of the eurocrate. The circuit only uses + 15V and + 24V, and I checked for continuity at all the IC pins marked with these voltages on the schematic.

Since the whole point of this test was to see if the slot I identified was delivering the right voltages, I think this is sufficient. I will now need to fashion a cable that I can use to connect a DC power supply to the PZT driver boards so that these can be tested further.

The high voltage points (100V DC) remain to be tested.

  8803   Thu Jul 4 19:37:37 2013 KojiUpdateSUSNew SUS screen

Totally agree. The old suspension screen should be driven away.

  8802   Thu Jul 4 17:14:53 2013 ranaUpdateSUSNew SUS screen

 Now that the 3f locking looks so cool for the PRMI, I suppose that the PRMI + arm stuff will be very successful.

At LLO, I've just noticed the screens that they have for the single pendulums / TTs. I'm attaching a screenshot of the one Zach is using for the steering into the OMC. We should grab these and replace our existing SUS screens with them.

Attachment 1: OM1.png
OM1.png
  8801   Thu Jul 4 01:09:33 2013 Annalisa, KojiUpdateLSCPRMI locked with REFL165

Summary:

- The new REFL165 PD was installed on the AP table
- The REFL165I/Q signals are now showing sensible and robust PRCL/MICH signals
- PRMIsb was locked only with these REFL165 signals

Details:

- Installation of the REFL165 PD

We prepared the REFL165 PD for the 4" optical height. The actual issue was the power supply for the PD.
We soldered wires between the PD and the RF PD interface break-out board. Then the PD interface
cable for the old REFL165 (iLIGO style) was connected.

At the REFL port, most of the light is rejected by the first beam splitter (R=90%?). We attenuated the beam by a factor of 10
using an ND filter. The new PD showed the DC output of ~10V. This corresponds to the photocurrent of 5mA.
(cf. the shot-noise intercept current is ~1mA)

The output of the REFL165 PD was checked with the RF spectrum analyzer. It was a bit surprising but we had a forest of
RF signals betwen 11MHz and 178MHz.
We tried to use a high-pass filter with fc=100MHz (SPH-100) but still the rejection
was not enough. We ended up with using SPH-150 (fc=150MHz).

- Whitening / Demodulation phase

Then we connected the RF output to the SMA cable to the LSC rack. We immediately saw the nice signals from REFL165I/Q
channels, namely sensible structure of pendulum resonances (1/3/16Hz peaks) and floor level.

The whitening level was changed from 21dB to 45dB (max). The DC offsets in the I/Q channels (of the order of 2000~4000)
were removed by the ./LSC/LSCoffset script.

Firstly we locked the PRMI with the usual signals (REFL33I and AS55Q).
The demodulation phase was roughtly tuned (1deg precision) such that the Q phase signal is minimized,
assuming most of the signal is coming from PRCL. Our choise is 74deg.

In this configuration, PRCL shows same quality of signal as our prefered PRCL (i.e. REFL33I) in the amplitude and the sign.

- Locking

We switched to the REFL165 signal by handing off at the input matrix. The input matrix element for REFL165_I was gradually
increasded up to 0.8 while the element for REFL33I was gradually reduced to 0. We did the same for REFL165_Q with the element of 0.2.

Now we tried locking with REFL165I/Q from the beginning. Once the alignment is adjusted, the lock was immediately obtained
only with REFL165I/Q.
Today we did not adjusted the ASC stuff (OPLEVs and PRM ASC) so the lock was not long (<1min). Particularly
ITMX poiting kept drifting and it made the lock difficult. We should check the oplev setup carefully.

- LSC summary

PRCL
Signal source: REFL165I (74deg) / Whitening gain 45dB
Normalization sqrt(POP110I x 0.1) / Trigger POP110I 100up 3down
Servo: input matrix 0.80 -> PRCL Servo FM3/4/5 Always ON G=+2.50
Actuator: output matrix 1.00 -> PRM

MICH
Signal source: REFL165Q (74deg) / Whitening gain 45dB
Normalization sqrt(POP110I x 10.0) / Trigger POP110I 100up 3down
Servo: input matrix 0.20 -> MICH Servo FM4/5 Always On G=-40
Actuator output matrix -1.00 -> ITMX / +1.00 -> ITMY

To Do:

- Refine the PRM asc servo (AC coupled)
- Align oplevs
- ITMX oplev is drifting quickly (~1min time scale)

 

 

 

 

Attachment 1: 130703_PRMI.pdf
130703_PRMI.pdf
Attachment 2: Screenshot.png
Screenshot.png
  8800   Wed Jul 3 21:19:04 2013 gautamConfigurationendtable upgradeplan of action for PZT installation

 This is an update on the situation as far as PZT installation is concerned. I measured the required cable (PZT driver board to PZT) lengths for the X and Y ends as well as the PSL table once again, with the help of a 3m long BNC cable, just to make sure we had the lengths right. The quoted cable lengths include a meter tolerance. The PZTs themselves have cable lengths of 1.5m, though I have assumed that this will be used on the tables themselves. The inventory status is as follows.

  1. Stuff ordered:
    • RG316 LEMO 00 (female) to SMB (female) cables, 10 meters - 6pcs (for the Y-end)
    • RG316 LEMO 00 (female) to SMB (female) cables, 11 meters - 6pcs (for the X-end)
    • RG316 LEMO 00 (female) to SMB (female) cables, 15 meters - 8pcs (6 for the PSL, and two spares)
    • RG316 SMA (male) to open cables, 3 meters - 3pcs (1 each for the X end, Y end and PSL table, for connecting the driver boards to the 100V DC power supply)
    • 10 pin IDC connectors for connecting the DAC interface to the PZT driver boards 
  2. Stuff we have:
    • 40 pin IDC connectors which connect to the DAC interface
    • PZT driver boards
    • PZT mounts
    • Twisted ribbon wire, which will be used to make the custom ribbon to connect the 10 pin IDC to the 40 pin IDC connector

I also did a preliminary check on the driver boards, mainly to check for continuity. Some minor modifications have been made to this board from the schematic shown here (using jumper wires soldered on the top-side of the PCB). I will have to do a more comprehensive check to make sure the board as such is functioning as we expect it to. The plan for this is to first check the board without the high-voltage power supply (using an expansion card to hook it up to a eurocrate). Once it has been verified that the board is getting powered, I will connect the high-voltage supply and a test PZT to the board to do both a check of the board as well as a preliminary calibration of the PZTs.

To this end, I need something to track the spot position as I apply varying voltage to the PZT. QPDs are an option, the alternative being some PSDs I found. The problem with the latter is that the interfaces to the PSD (there are 3) all seem to be damaged (according to the labels on two of them). I tried connecting a PSD to the third interface (OT301 Precision Position Sensing Amplifier), and hooked it up to an oscilloscope. I then shone a laser pointer on the psd, and moved it around a little to see if the signals on the oscilloscope made sense. They didn't on this first try, though this may be because the sensing amplifier is not calibrated. I will try this again. If I can get one of the PSDs to work, mount it on a test optical table and calibrate it. The plan is then to use this PSD to track the position of the reflected beam off a mirror mounted on a PZT (temporarily, using double sided tape) that is driven by feeding small-amplitude signals to the driver board via a function generator. 

 

Misc

The LEMO connector on the PZTs have the part number LEMO.FFS.00, while the male SMB connectors on the board have the part number PE4177 (Pasternack)

Plan of Action:

  • The first task will be to verify that the board is working by the methods outlined above.
  • Once the board has been verified, the next task will be to calibrate a PZT using it. I have to first identify a suitable way of tracking the beam position (QPD or PSD?)
  • I have identified a position in the eurocrate at 1Y4 to install the board, and I have made sure that for this slot, the rear of the eurocrate is not hooked up to the cross-connects. I now need to figure out the exact pin configuration at the DAC interface: the bank is marked 'DAC Channels 9-16' (image attached) but there are 40 pins in the connector, so I need to map these pins to DAC channels, so that when making the custom ribbon, I get the pin-to-pin map right.

DAC_bank.png

 

The wiring scheme has been modified a little, I am uploading an updated one here. In the earlier version, I had mistaken the monitor channels as points from which to log data, while they are really just for debugging. I have also revised the coaxial cable type used (RG316 as opposed to RG174) and the SMB connector (female rather than male).

ASC_schematic.pdf 

 

 

 

 

  8799   Wed Jul 3 20:51:43 2013 CharlesUpdateISSProposed ISS for CTN Experiment - Altium Schematic

 After familiarizing myself with Altium, I drew up the attached schematic for the ISS to be used in the CTN experiment. The filename includes 'abbott-switch' as I am using an Altium component (the switch, in particular), that he created. The MAX333A actually has 20 pins on a single component, but the distributed component that he created is useful for drawing uncluttered schematics. I won't be using all of the pins on this switch, but for completeness, I have included the 3rd and 4th portion of the full component in the upper right hand corner.

Currently, the schematic includes the voltage reference (AD586), a LP filter for the reference signal, the differential amplifier stage to obtain the error signal and then finally all of the filter stages. The schematic does not include the RMS detection and subsequent triggering of each filter stage. The TRIGGER 1 signal is a user input (essentially the on button) while the TRIGGER 2 signal will flip the second switch when the RMS noise has decreased sufficiently after the first filter stage has been turned on. 

PCB layouts will be done once I understand that part of Altium 

 

NOTE THAT I HAVE DELETED ELOG 8798 AS IT WAS A DUPLICATE OF THIS ONE.

I wanted this elog to be in reply to a previous one and I couldn't figure out how to change that in an elog I already submitted.

 

 

 

Attachment 1: CTNServo_v2_abbott-switch.pdf
CTNServo_v2_abbott-switch.pdf
  8797   Wed Jul 3 14:33:46 2013 KojiSummaryLSCTest result for the REFL165 photodetector

P.1 Circuit diagram

Added components are indicated by red symbols.

- The diode on the board is HAMAMATSU S3399. It is a Si PIN diode with φ3.0 mm.

- Based on prototype version of aLIGO BBPD D1002969-v8 (although the board says v7, It is v8.)

- The input impedance of the MAR-6SM amplifier (50Ohm) provides the transimpedance.

- The first notch (Lres and Cresa/b) is actually not notch but a LF rejection with DC block.

- The second and third notches are tuned to 11MHz and 55MHz.

- Another notch is implemented between the RF amps. The 33MHz signal is weak so I expected
to have no saturation at the first amplifier.

- As you see from the DC path, the transimpedance of the DC path is 2k V/A. If this is too high,
  we need to replace R9 and R11 at the same time. TP1 is providing +10V such that the total
  reverse bias becomes 25V without bringing a special power supply.

P.2 Transimpedance

The transimpedance is measured with an amplitude modulated diode laser.

The transimpedance is 1k V/A ish. It is already at the edge of the bandwidth.
If we need more transimpedance at 165MHz, we should replace
the PD with FFD-100 (I have one) and apply 100V of reverse bias.

P.3 Current noise spectrum

The measured dark noise voltage spectrum was converted to the equivalent current noise at the diode.

The measured transimpedance is ~1.2kV/A.
The reduction of the transimpedance above 100MHz has been seen as 165MHz is already at the edge of the bandwidth.
If we need more transimpedance at 165MHz, we should replace the diode with FFD-100 (I have one) and apply 100V of reverse bias.

P.4 Shot-noise intercept current

Shot-noise intercept current was measured with a white light from a light bulb.
This measurement suggests the shot-noise intercept current of 1mA, and transimpedance of 1.5kV/A.

Attachment 1: REFL165_response_130702.pdf
REFL165_response_130702.pdf REFL165_response_130702.pdf REFL165_response_130702.pdf REFL165_response_130702.pdf
  8795   Wed Jul 3 11:07:17 2013 AlexSummary Photodetector Characterization

 [Alex, Koji]

We characterized Koji's BBPD MOD for REFL165 (see attachment).

First, we calibrated the Agilent 4395 Network Analyzer (NA) to account for differences in cable features between the Ref PD and Test PD connections. This was done using the 'Cal' softkey on the NA. 

Then we performed transimpedance measurements for the test PD and reference PD relative to the RF output of the NA and relative to each other (see 2nd attachment. Note that the NA's RF output is split and sent to both the IR Laser and the NA's Ref input).

Next, we made DC measurements of the outputs of the photodetectors to estimate the photocurrent distribution of the transimpedance setup (like the 2nd attachment, but with the outputs of the PDs going to a multimeter). By photocurrent distribution, we mean how the beamsplitter and respective quantum efficiencies/generalized impedance/etc. of the PDs influence how much current flows through each PD at with a DC input.

Finally, we measured the output noise as a function of photocurrent (like the 2nd attachment, but with a lightbulb instead of the IR Laser). Input voltages for the lightbulb ranged from 0mV to 6V. Data was downloaded from the NA using netgpibdata from the scripts directory. Analysis is currently in progress; graphs to come soon.

 

Attachment 1: BBPD_PCB.pdf
BBPD_PCB.pdf
Attachment 2: transimpedance_measurement.pdf
transimpedance_measurement.pdf
  8794   Wed Jul 3 10:39:25 2013 manasaUpdateIOOMC aligned and WFS enabled

I found WFS had been left disabled from sometime yesterday. I don't see anyone mentioning  when and why they had turned OFF the WFS servo.

I aligned MC and turned ON the WFS servo. MC is back.

  8793   Wed Jul 3 03:06:29 2013 AnnalisaUpdateGreen LockingALS servo configuration

 

I realized that I cannot open the attached plots. I'll fix them tomorrow.

  8792   Wed Jul 3 01:49:46 2013 AnnalisaUpdateGreen LockingALS servo configuration

[Koji, Annalisa, Manasa]

Today we worked on the ALS servo stabilization for the Y arm.

First step: find the beat note

The beat note was found following the usual steps:

  • Y arm cavity locked on IR to have a good alignment
  • Y arm cavity locked on green (eventually unlocked on IR)
  • beat note alignment maximized on the PSL table

Beat note amplitude = -27 dBm @ 50 MHz

PSL temperature = 31.54 degC

Laser Offset on the slow servo2 = -11011

 

In the GREEN HORNET we did the following changes for the Y arm:

Input Signal Conditioning

On the C1ALS-BEATY_FINE  screen the same antiwhitening filters of the C1ALS-BEATX_FINE have been reproduced. At moment, only the FM3 [10:1] is enabled.

On the C1ALS-BEATY_FINE_PHASE screen the gain was set at 3600, since the amplitude of the Q signal after the Phase rotator (BEATY_FINE_Q_ERR) was about 30. To set this value we made a proportion with respect to a previous optimized value, where the amplitude was 100 and the gain was set to 1200.

DOF filters

In order to stabilize the beat frequency, we started enabling the FM5 [1000:1] filter in the C1ALS_YARM panel, and then we started increasing the gain first in small steps (0.1), in order to understand which sign the gain should have without kicking the mirror.

We measured the Power Spectrum of the C1:ALS-BEATY_FINE_PHASE_OUT in-loop signal while varying the gain of the C1ALS_YARM servo filter.

Eventually, we enabled the following filters:

FM2 [0:1]

FM3 [1:5]

FM4 [1:50]

FM5 [1000:1]

FM6 [RG3.2]

FM7 [RG16.5]

Gain = -30.

Koji expects the UGF of the loop to be around 100-ish Hz, and he also expected the small bump around 300-400 Hz.

Then we realized that the channel we were measuring was not calibrated in unit of Hz, so we took again the measurement looking at the channel C1:ALS-BEATY_FINE_PHASE_OUT_HZ. In this case, we didn't observe any bump. Maybe the beat frequency was slightly changed from the previous measurement and the all servo shape was also different. The final value of the gain was set at -8.

The Y axis unit is missing (bad me!). It's in deg/sqrt(Hz) for the first plot and Hz/sqrt(Hz) for the second one.

 

Attachment 1: ALS
Attachment 2: ALS_calibrated
  8791   Tue Jul 2 12:59:46 2013 CharlesUpdateISSGeneral Design for ISS Applicable to Multiple Applications

 While attempting to develop a somewhat accurate noise budget for the 40m, I reasoned that while the shape of the transfer function for the ISS is important, the degree to which we can 'tune' it to a particular experiment/application is limited.

  • Since we're using a DC-coupled servo, the TF magnitude will go like f^k with k < 0 for low frequency.
  • The UGF will be somewhere around 10 kHz to 1 MHz (most likely right around 100 kHz) as beyond 1 MHz, the gain of our servo is limited by the GBWP of the op-amps.
  • We need around 3 or 4 orders of magnitude of gain in the 1-100 Hz range based on this, with gain > 10 for f < 10 kHz

Beyond that, we're sort of limited by the desired high and low frequency behavior as well as the general principle that more electronics = more noise so we probably don't want more than 3 or 4 filter stages, if that. Additionally, the ISS can be over-engineered so that it suppresses the laser noise to levels well below other fundamental noise sources over the important regime ~10 - 10^3 Hz without particular regard to a noise budget.

The design I propose is very similar to a previous design, with a few adjustments. It consists of 3 filter stages that easily be modified to increase gain for higher frequencies if it is known/determined that the laser being stabilized has a lot of high frequency noise.

40mServo_v1.png

Stage 1: Basic LP Filter + Establish UGF (each stage 'turning on' will not change the UGF),  Stage 2: Integrator with zero @ 10 kHz,  Stage 3: Optional extra gain if necessary

40mServo_v1-Stage1.pdf40mServo_v1-Stage2.pdf40mServo_v1-Stage3.pdf

With the full TF given by,

 40mServo_v1.pdf 

As usual we consider the noise caused by the servo itself. Noise analysis in LISO is done with a 1 V input excitation.

40mServo_v1-Input_Noise.pdf

This servo should function sufficiently for the 40m.

  8790   Tue Jul 2 02:29:47 2013 JenneUpdateLSCSensing Matrix vs. Xarm sweep

Here is the Sensing Matrix movie (sorry for the iffy quality - my movies usually come out better than this):

 

This is the sensing matrix for the sideband locked on PRMI, bringing the Xarm into resonance from anti-resonance, in 20 equally-spaced steps.  You can see the microscopic ETMX offset (units of meters) in the title of the figures.

I was surprised to see some of the 'jumps' in the sensing matrix that happen near the end, when the arm is almost in resonance.  I'm in the process of making movies of the error signals as the Xarm is brought into resonance.  I'll have to post those in the morning, since they're taking a long time to produce and save, however when I looked at a few, there is some weird stuff going on as we get close to resonance, even with the 3f signals. 

The modeling phone call is in the morning, but if anyone who is not regularly on the call has thoughts, I'm all ears.

  8789   Tue Jul 2 00:25:14 2013 gautamUpdateGreen LockingUniversal PDH box tuning

 [Koji, Annalisa, Gautam]

Annalisa noticed that over the weekend the Y-arm green PDH was locked to a sideband, despite not having changed anything on the PDH box (the sign switch was left as it was). On friday, we tried turning on and off some of the filters on the slow servo (C1ALS_Y_SLOW) which may have changed something but this warranted further investigation. We initially thought that the demodulation phase was not at the optimal value, and decided to try introducing some capacitances in the path from the function generator to the LO input on the universal PDH box. We modelled the circuit and determined that significant phase change was introduced by capacitances between 1nF and 100nF, so we picked out some capacitors (WIMA FKP) and set up a breadboard on which to try these out.

After some trial and error, Koji dropped by and felt that the loop was optimized for the old laser, the various loop parameters had not been tweaked since the new laser was installed. The following parameters had to be optimized for the new laser;

  • Servo gain
  • LO frequency
  • LO modulation depth
  • Demodulation phase

The setup was as follows: 

  • PDH box error signal to Oscilloscope CH1
  • Green PD output to Oscilloscope CH2
  • No capacitor between Function Generator and the PDH box
  • 0.1Hz triangle wave (30 counts amplitude) applied to ETMY via awggui (so as to sweep the cavity and see stronger, more regular TEM00 flashes)

The PDH error signal did not have very well-defined features, so Koji tweaked the LO frequency and the modulation depth till we got a reasonably well-defined PDH signal. Then we turned the excitation off and locked the cavity to green. The servo gain was then optimized by reducing oscillations in the error signal. Eventually, we settled on values for the Servo Gain, LO frequency and modulation depth such that the UGF was ~20kHz (determined by looking at the frequency of oscillation of the error signal on an Oscilloscope), and the PDH signal had well-defined features (while the cavity was unlocked). The current parameters are

  • LO frequency: 205.020 kHz
  • modulation depth: 0.032 Vpp

We then proceeded to find the optimal demodulation phase by simulating the circuit with various capacitances between the function generator and the PDH box (circuit diagram and plots attached). The simulation seemed to suggest that there was no need to introduce any additional capacitance in this path (introducing a 1nF capacitance added a phase-lag of ~90 degrees-this was confirmed as the error-signal amplitude decreased drastically when we hooked up a 1nF capacitor on our makeshift breadboard). In the current configuration, the LO is connected directly to the PDH box.

 

Misc Points:

  • The phase shifter in the PDH box is not connected: the IC in the box, JSPHS-26, is designed for operation in the range 18-26MHz. If necessary, an all-pass-filter could be introduced, with a tuneable rheostat to adjust the phase for our frequency range. Right now, turning the knob marked "LO phase angle" on the front panel doesn't do anything. The mixer on the PDH board is also not used for the same reason.
  • PSL shutter was closed sometime earlier this evening, because we suspected some IR light was reaching the Green PD on the y-endtable, and was influencing the error signal. Its back open now.
  • Useful information about the old y-end laser relevant to selecting the right LO frequency, modulation depth, and servo gain can be found here and in elog 2746 and subsequent replies, though the details of how the measurement were made aren't entirely clear. The idea is that the characteristics of the piezoelectric element in the laser has some characteristics which will determine the optimal LO frequency, modulation depth and servo gain.

 

To Do:

Now that we are reasonably confident that the loop parameters are optimal, we need to stabilise the C1ALS_Y_SLOW loop to stabilise the beat note itself. Appropriate filters need to be added to this servo.

 

Circuit Diagram: 50 ohm input impedance on the source, 50 ohm output impedance seen on the PDH box, capacitance varied between 1nF and 100nF in steps.

circuit.pdf

Plots for various capacitances: Gold-green trace (largest amplitude) direct from LO, other traces at input to PDH box.

model.pdf

  8788   Mon Jul 1 23:27:07 2013 JenneUpdateLSCSensing Matrix vs. Xarm sweep

I have modeled the PRMI sensing matrix as I bring the Xarm into resonance.  In optickle, I have the PRMI on sideband resonance, the ETMY is artificially set to have a transmission of 1, and the ETMX has it's nominal transmission of 15ppm.  I start with the ETMX's microscopic position set to lambda/4 (antiresonant for IR in the arm), and take several steps until the ETMX's microscopic position is 0 (resonant for IR in the arm).

Xarm antiresonant:

Modeled sensing matrix, units = W/m, Offset = 2.66e-07, phase in degrees
 
            MICH Mag   MICH Phase    PRCL Mag   PRCL Phase  
AS55         3.348E+04   142.248      5.111E+03    70.571    
POX11        3.968E+01   -66.492      1.215E+04    54.312    
REFL11       3.231E+05    24.309      9.829E+07   144.311    
REFL165      9.946E+03  -159.540      4.540E+05   -64.710    
REFL33       1.963E+04  -168.530      1.573E+06    -2.744    
REFL55       1.160E+06    -6.755      5.429E+07    86.895 

 

Xarm resonant:

Modeled sensing matrix, units = W/m, Offset = 0, phase in degrees
 
            MICH Mag   MICH Phase    PRCL Mag   PRCL Phase  
AS55         1.647E+06    57.353      3.676E+06   -81.916    
POX11        3.927E+02  -118.791      2.578E+04  -102.158    
REFL11       7.035E+05    61.203      1.039E+08   167.149    
REFL165      1.602E+04  -144.586      5.971E+05   -49.802    
REFL33       2.157E+04   171.658      1.940E+06    -9.133    
REFL55       1.822E+06     7.762      6.900E+07   101.906 

 

For REFL55, the MICH magnitude increases by a factor of 1.6, while the PRCL  magnitude increases by 1.3 .  The MICH phase changes by 15 degrees, while the PRCL phase also changes by 15 degrees.  Just eye-balling (rather than calculating), the other REFL PDs look to have similar-ish magnitude and phase changes.  Certainly none of them are different by orders of magnitude.

Movies forthcoming.

  8787   Fri Jun 28 17:33:33 2013 John ZweizigUpdateSUSNDS2 Status

Quote:

Quote:

I've restarted the NDS2 process on Megatron so that we can use it for getting past data and eventually from outside the 40m.

1) from /home/controls/nds2 (which is not a good place for programs to run) I ran nds2-megatron/start-nds2

2) this is just a script that runs the binary from /usr/bin/ and then leaves a log file in ~/nds2/log/

3) I tested with DTT that I could access megatron:31200 and get data that way.

There is a script in usr/bin called nds2_nightly which seems to be the thing we should run by cron to get the channel list to get updated, but I' m not sure. Let's see if we can get an ELOG entry about how this works.

Then we want Jamie to allow some kind of tunneling so that the 40m data can be accessed from outside, etc.

 I have done the following:

  * installed the nds2-client in /ligo/apps/nds2-client

  * moved the nds2 configuration directories to /ligo/apps/nds2/nds2-megatron

  * set up a cron job to update the channel list every morning at 5 am. The cron line is:

     15 5 * * * /usr/bin/nds2_nightly /ligo/apps/nds2/channel-tracker /ligo/apps/nds2/nds2-megatron

    cron will send an email each time the channel list changes, at which point you will have to restart the server with:

     cd /ligo/apps/nds2/nds2-megatron
     pkill nds2
     ./start-nds2

  * restarted nds2 with updated channel lists.

 I have set the cron job up to restart the nds2 server automatically if the channel list changes. The only change is that the cron command was changes to /ligo/apps/nds2/nds2-megatron/test-restart.

 

  8786   Fri Jun 28 16:19:06 2013 CharlesUpdateISS40m Noise Budget - Seismic Contribution

 I'm working on developing a full noise budget for the 40m. To that end, I'll use measurements from the GUR1 seismometer to characterize seismic noise. Without any unit calibration, I found the following spectrum,

seismic_noise_6-28-13_raw.png

To extract useful information from this data, I first used the calibration from "/users/Templates/Seismic-Spectra_121213.xml" to obtain the spectrum in [m / s / sqrt(Hz)].

calibrated_data = raw_data * 3.8e-09

I then divided each point in the power spectrum by the frequency of said point to obtain [m / sqrt(Hz)]. I don't think we can simply divide the whole spectrum by 40 meters to obtain [RIN / sqrt(Hz)], although that was my immediate intuition. Having power spectra of all the major noise contributions in units of [RIN / sqrt(Hz)] would make designing an appropriate filtering servo fairly straightforward.

 seismic_noise_6-28-13_meters.png

 

Attachment 2: seismic_noise_6-28-13_raw.fig
Attachment 4: seismic_noise_6-28-13_meters.fig
  8785   Fri Jun 28 14:57:15 2013 SteveUpdateGeneralclosing lab doors

I found the south end emergency doors not latched completely. There was a ~ 3/8" vertical gap from top to bottom.

Please pull or push  doors harder if they not catch fully.

  8784   Fri Jun 28 13:10:28 2013 gautamUpdateGeneralITMx Oplev-servo gains adjusted

 

 As mentioned in elog 8770, I wanted to give the POX beam a little more clearance from the pick-off mirror steering the outcoming oplev beam. I tweaked the position of this mirror a little this morning, re-centred the spot, and checked the loop transfer function once again. These were really close to those I measured last night (UGF for pitch ~8Hz, for yaw ~7Hz), reported in elog 8777, so I did not have to change the loop gains for either pitch or yaw. Plots attached.

 pitch-plot_copy.pdf

 

yaw-plot_copy.pdf

  8783   Fri Jun 28 12:15:09 2013 KojiUpdateLSCNeed to measure sensing matrix at REFL165

There is no sensible REFL165 PD in the lab. I am supposed to prepare a new version of REFL165 using prototype BBPD.

ELOG V3.1.3-